I'm having trouble factoring out this problem, I didn't know where to post a factoring question so I apologize if this is the wrong place, but im trying to factor out x^2 - 15x - 2 I tried isolating the coefficient and constant and cross multiply to get the two numbers that = -15x, but I'm stumped lol, I'd appreciate a hand! Thannks!
Answer by mananth(16949)

x^2 - 15x - 2
This can be done in many ways
x^2-15x+(7.5)^2-2=0
x^2-15x+(7.5)^2=2+(7.5)^2
(x-7.5)^2= 2+56.25
(x-7.5)^2=58.25
Take the square root of both sides
(x-7.5)= +/- sqrt(58.25)
(x-7.5)= 7.63
x= 7.5+7.63= 15.13
OR
x= 7.5-7.63= -0.13
m.ananth@hotmail.ca